# Hidden in Plain Sight II: Characterizing the luminous companion to $\kappa$ Velorum with VLTI/GRAVITY

### v1

@@ Abstract
$\kappa$ Vel (Markeb, HD 81188) is one of the brightest stars in the Southern sky and has long been known to be a single-lined spectroscopic binary. The binary mass function is large, $f(M)=1.15\ M_{\odot}$, suggesting that the bright ($V=2.5$) B2IV star may host a dark, compact object companion. We use VLTI GRAVITY observations to definitively test this possibility by directly resolving the binary. We detect a main sequence B star companion and rule out the compact object scenario. By combining the relative astrometric orbit and archival radial velocities, we report an updated precise characterization of the orbit (period $P=116.795\pm 0.002$ d, eccentricity $e=0.1764\pm 0.0004$, inclination $i=74.04\pm 0.01^{\circ}$) and estimate the masses of the B stars. Using the original Hipparcos parallax measurement $\varpi=6.05\pm 0.48$ mas, we find $M_{1}=10^{+4}_{-2}\ M_{\odot}$ and $M_{2}=6.9\pm 1.0\ M_{\odot}$. The uncertainties on the masses are primarily driven by the uncertain parallax, which we find is likely biased by the orbital motion. We use an archival UVES spectrum and MIST evolutionary tracks to refine our mass estimates. Finally, we discuss how interferometry and high-contrast imaging may be used to characterize other candidate star+compact object binaries, including those that will be discovered with Gaia DR4, as part of a larger effort to uncover the hidden population of black holes in the Milky Way.
@@ 1 Introduction
The majority of the observed black holes and neutron stars in the Milky Way have been detected in X-ray binary systems (CorralSantana16). While compact object masses can sometimes be measured dynamically in these systems, either by inferring the orbital inclination from the accretion disk properties (e.g., Torres20) or by modeling ellipsoidal modulations (e.g., Orosz01), many of the transient BH candidates have no mass measurements (e.g., Russell22). Since X-ray binaries are expected to undergo at least one episode of binary interaction and mass transfer (Kalogera96; Tauris06), the observed compact object mass distribution from X-ray binaries may present a biased view of the intrinsic distribution.
@@
Outside of the Galaxy, gravitational waves have been used to detect >150>150 compact object binary mergers (Ligo2025). While these targets provide powerful constraints on the mass distribution of black holes (Ligo2025_population), these rare extragalactic systems are also the products of complex binary interactions that result in a coalescence. The majority of the observed gravitational wave sources are also more massive than typical X-ray binaries, since more massive black hole mergers have a larger characteristic strain.
@@
In recent years, there has been an increased effort to uncover the hidden population of dormant black holes in the Milky Way. Isolated black holes can only be detected via microlensing (Lam22; Lam23). While there is currently only one strong candidate, the Vera C. Rubin Legacy Survey of Space and Time and the Roman Space Telescope are expected to detect $\gtrsim 30$ isolated black holes (Abrams25; Lam23_whitepaper). Multiple observational methods can be used to detect and characterize black holes in wide binaries with luminous companions. Three black holes have been detected using Gaia astrometry (ElBadry23_BH1; ElBadry23_BH2; Chakrabarti23; Tanikawa23), including Gaia-BH3, a red giant star in the Milky Way halo with a $33\ M_{\odot}$ black hole in an 11.9 year binary orbit (GaiaBH3). Spectroscopic surveys have also been applied to search for candidates (e.g., Giesers18; Shenar22; Mahy22), but the unknown orbital inclination means that only a minimum black hole mass can be measured. At short orbital periods ($P\lesssim 10$ days for main sequence companions), ellipsoidal modulations can be used to measure the orbital inclination (Morris93), but no strong candidates have been discovered starting from photometric variability (Rowan21; Green25).
@@
Regardless of whether a black hole candidate is detected with astrometry, spectroscopy, or photometry, confirming a black hole effectively means ruling out any type of luminous companion. For some systems, like the $33\ M_{\odot}$ black hole in Gaia BH3, the spectral energy distribution can easily exclude a massive main sequence star or an equal-mass binary of two $16.5\ M_{\odot}$ stars. For lower-mass black holes, spectral disentangling can be help to distinguish compact objects from false positives, such as rapidly rotating subgiants (e.g., ElBadry22_zoo; Jayasinghe22). However, if we can directly resolve a candidate system with high-contrast imaging or interferometry, we can detect or rule out luminous companions with a single observation (e.g., Frost22).
@@
If we observe a stellar binary, the direct imaging observations will resolve the two stars and help constrain the three-dimensional orbit. If we instead resolve a non-interacting black hole binary, we expect to observe a single point source. The combination of RVs and astrometry can therefore be used to unambiguously detect and characterize non-interacting black hole binary systems. In Rowan25_shark, we performed a pilot study to explore the feasibility of this method using SHARK-VIS (Pedichini22; Pedichini24) on the Large Binocular Telescope (Hill10). We ruled out black holes for three of the four systems observed, and showed that two of the binaries selected are likely hierarchical triples. Since the SHARK-VIS observations only require $\sim 10$ min of on-sky observing time, direct imaging can be used to rapidly vet black hole candidates in long period orbits more efficiently than high-resolution spectroscopic follow-up.
@@
Here, we use VLTI/GRAVITY to observe $\kappa$ Velorum (hereafter $\kappa$ Vel), a bright ($V=2.5$), single-lined spectroscopic binary with an RV orbit that indicates a large, massive companion. Section §2 describes how we selected this target as a potential black hole candidate. Section §3 reports the VLTI/GRAVITY observations and the describes the data reduction. We unambiguously detect a second star in the VLTI GRAVITY observations and rule out a black hole. We then characterize the stellar binary by simultaneously fitting the archival RVs with the relative astrometry in Section §4, model an archival high resolution spectrum (Section §5), and make comparisons to evolutionary tracks of massive stars (Section §6). Finally, we discuss the prospects for observing other high-mass function single-lined spectroscopic binaries with interferometry and direct imaging in Section §7.
2 Target Selection
@@
We selected targets following the procedure outlined in Rowan25_shark. In short, we started from two large catalogs of spectroscopic binaries, the Ninth Catalog of Spectroscopic Binary Orbits (Pourbaix04, hereafter SB9) and single-lined spectroscopic binaries from the Gaia DR3 non-single stars catalog (GaiaHiddenTreasure; Gosset24, hereafter Gaia SB1). We used the astrophysical parameter estimates from StarHorse (Anders22) to estimate the mass of the photometric primaries, and determined the minimum companion masses and maximum projected orbital separations based on the archival RV orbits. Figure 1 shows the maximum projected orbital separations and minimum companion mass measurements.
@@
We identified $\kappa$ Vel (HD 81188, HIP 45941) as having a large minimum companion mass and an angular separation accessible with VLTI GRAVITY. With an apparent magnitude of $V=2.5$, $\kappa$ Vel is one of the brightest stars in the Southern sky and has been targeted as part of a number of surveys, but the secondary has not previously been characterized. $\kappa$ Vel has been most commonly observed as part of interstellar medium absorption line studies (e.g., Crawford91; Fruscione94; Jenniskens97; Crawford02; Smith13; Cox17). $\kappa$ Vel was first identified as a binary candidate by Wright05 and the orbit was characterized by Curtis1907 as part of the D.O. Mills Expedition from Lick Observatory to the Southern Hemisphere. Five additional measurements taken by Buscombe60 confirmed the RV variability and orbit. Table 1 lists the archival measurements. We assume an RV uncertainty of 5 km/s for the Curtis1907 measurements since measurement errors were not reported. Chini12 also measured four RVs of $\kappa$ Vel as part of a spectroscopic surveys of massive stars. They classify the system as an SB1, but the individual RV measurements or spectra are not available. $\kappa$ Vel is included in the SB9 catalog with an orbital period of $P=116.65$ days, $K_{1}=46.5$ km/s, and a binary mass function $f(M)=1.15\ M_{\odot}$, where
@@
$M_{1}$ is the photometric primary mass, $M_{2}$ is the companion mass, $i$ is the orbital inclination, and $e$ is the orbital eccentricity. The photometric primary is spectroscopically characterized as a B2IV star (Levato75), which roughly corresponds to a primary mass of $\sim 7$$10\ M_{\odot}$ (Pecaut13). Based on the RV orbit, even if the primary is $7\ M_{\odot}$ and the orbit is edge-on, the companion must be $5.7\ M_{\odot}$ (Figure 2). Because of the large minimum companion mass, Trimble69 included $\kappa$ Vel in their catalog of binaries that could host compact object companions. The companion could be another massive star, a compact object, or an inner binary of two lower mass stars. The maximum projected orbital separation is >7.0>7.0 mas (Figure 1), making this a suitable target for VLTI direct imaging observations to test these possibilities.
Figure 1: Minimum companion masses, $M_{2,\rm{min}}$, and maximum projected angular separation for binaries in the Gaia SB1s and SB9 catalogs. $\kappa$ Vel is marked in blue. The vertical line marks the projected angular separation resolvable with VLTI GRAVITY.
Table 1: Archival RV measurements of $\kappa$ Vel from Curtis1907 and Buscombe60. Since Curtis1907 do not report RV uncertainties, we estimate a conservative RV uncertainty of 5 km/s. We also include our RV measurement from the UVES spectrum. We assign an RV uncertainty on the UVES spectrum of 22 km/s to account for the unknown systematic RV offset between the archival measurements and the UVES measurement.
@@
3 VLTI/Gravity Observations
@@
Table 2: VLTI GRAVITY observation log. The dates are given for the start of the night. The calibrator uniform-disk (UD) diameters are estimated with the JMMC SearchCal tool (Chelli16).
@@
3.1 Astrometry model-fitting
The visibility and closure phase data shows strong sinusoidal modulations that are indicative of a close binary system. In order to quantify the binary star parameters, we fit geometric models with the PMOIRED modeling software (Merand22). The free parameters are the relative separation vector of the two stars measured from the primary to the secondary (dRA, dDEC) and their $K$-band flux ratio ($\epsilon=f_{2}/f_{1}$). The stars are found to be unresolved, so we fixed the uniform-disk diameters in the fit to 0.1 mas. For the error estimation, we bootstrap over the individual measurements that were taken during each epoch. The derived astrometry is reported in Table 3.
Table 3: Astrometry for $\kappa$ Vel. The measurement uncertainties are estimated by bootstrapping over the individual measurements taken during each epoch (Section §3.1). $\rho$ is the correlation coefficient between the dRA and dDEC measurements and $f_{2}/f_{1}$ is the flux ratio in the $K$-band.
@@ 4 Orbit Fitting
We simultaneously model the RVs and relative astrometry with Markov Chain Monte Carlo (MCMC) with emcee (ForemanMackey13). We sample over six orbital parameters: the period, $P$, the periastron time, $T_{0}$, the orbital inclination, $i$, the longitude of the ascending node, $\Omega$, the eccentricity, $e$, and the longitude of periastron, $\omega$. We parameterize the latter two as $\sqrt{e}\cos\omega$ and $\sqrt{e}\sin\omega$ for computational efficiency. The angular size of the astrometric orbit is the ratio of the orbital semi-major axis to the distance, $a_{0}=a/d$. We sample over the velocity-semi amplitude of the photometric primary, $K_{1}$, and the system velocity, $\gamma$. We include an offset $\delta_{\rm{RV}}$ between the Curtis1907 and Buscombe60 RVs to account for systematic differences between the instruments. We use a Gaussian prior on $\delta_{\rm{RV}}$ centered on zero with $\sigma=1.5$ km/s. Since there is only one UVES spectrum, we do not fit for the potential RV offset between the Curtis1907 RVs and the UVES spectrograph, and instead choose to inflate the measured UVES RV uncertainty (see Section §5).
Finally, we include two additional terms to account for potentially underestimated uncertainties. The stellar jitter term, $s$, accounts for stellar variability and underestimated RV uncertainties (PriceWhelan17). To account for underestimated astrometric uncertainties, we also include a multiplicative term, $s_{\rm{ast}}$, that scales the bootstrapped astrometric uncertainties.
@@
where $\text{RV}_{\text{pred}}(t_{i})$ is calculated using the standard Keplerian orbit equations. The astrometric orbit term is
@@
$G$ $=a_{0}\left(-\sin\Omega\sin\omega+\cos i\,\cos\Omega\cos\omega\right)$
We use 20 walkers and run the chains for 500,000 iterations. We visually inspect the walker distributions and discard the first 100,000 iterations as burn-in. We find that the Gelman-Rubin statistic is $\hat{R}<1.01$ and the effective sample size is >10,000>10{,}000 for all parameters, suggesting that the chains have converged. Figure 4 shows the corner plot of the MCMC posteriors and random samples of the orbit model compared to the VLTI astrometry observations and Figure 3 shows the resulting orbit model. Table 4 reports the median and $1\sigma$ uncertainties of the posteriors.
Given a parallax, $\varpi$, and associated uncertainty, $\sigma_{\varpi}$, the binary masses $M_{1}$ and $M_{2}$ can be solved for using Kepler’s third law. Figure 5 shows the mass posteriors using the original Hipparcos parallax $\varpi=6.05\pm 0.48$ mas (Esa97_publication). The revised Hipparcos parallax, which primarily includes improvements to attitude reconstruction, is $\varpi=5.7\pm 0.3$ mas (vanLeeuwen07). The revised Hipparcos astrometry is a “stochastic” solution, rather than the standard five-parameter solution, which includes an additional “cosmic noise” astrometric noise term. Unfortunately, although $\kappa$ Vel has been observed by Gaia, since it is $G=2.5$ mag, it is likely saturated (Fabricius21) and has no reported parallax.
@@
$\alpha=a\left(\frac{q}{1+q}-\frac{\epsilon}{1+\epsilon}\right)$ (8)
@@
We must estimate the flux ratio of the system in the Hipparcos $H_{p}$ filter to determine the parallax bias. For a given mass pair of $M_{1}$ and $M_{2}$ and metallicity, we interpolate the MIST evolutionary tracks (Dotter16; Choi16) onto the same age grid, calculate the $K$-band flux as a function of stellar age, and use the observed flux ratio to estimate the system age. We use the $H_{p}$-band flux ratio at this system age to calculate $\alpha\sin i$. We find $\alpha\sin i\approx 0.22$, which indicates that the parallax can be biased by the binary motion by as much as 22%. Since the mass ratio depends on the parallax, this fractional bias does increase with increasing distance, but only at the percent level.
Figure 3: Top: RV orbit model and residuals. The red lines show random samples from the MCMC posteriors. Bottom left: VLTI GRAVITY astrometry measurements (black points) and random samples for the astrometric orbit model (blue). Bottom right: residuals for the astrometric orbit model. The solid errorbars give the uncertainty derived from the PMOIRED model, and the dashed errorbars report the scaled uncertainty $\sigma_{i}^{\prime}=s_{\text{ast}}\sigma_{i}$.
@@
5 UVES Spectra
$\kappa$ Vel was previously observed with VLT UVES as part of the ESO Diffuse Interstellar Bands Large Exploration Survey (EDIBLES, Cox17, 194.C-0833). $\kappa$ Vel was observed on 2017-04-18 in all four UVES cross-dispersers (CDs). The Phase 3 data product includes the combined spectra for 20 individual exposures in CD1 and CD2 and for 14 and 15 exposures in CD3 and CD4, respectively. While the purpose of the EDIBLES program was to characterize diffuse interstellar bands (DIBs), we can also use the UVES $\kappa$ Vel spectrum to characterize the photometric primary of the binary. We start by continuum normalizing the CD2 spectrum with a cubic spline, which covers a wavelength range of 370–500nm. We mask the Balmer lines during continuum normalization. Our analysis only uses the CD2 observations given the density of spectral features and lack of telluric lines.

### v2

@@ Abstract
$\kappa$ Vel (Markeb, HD 81188, 天社五) is one of the brightest stars in the Southern sky and has long been known to be a single-lined spectroscopic binary. The binary mass function is large, $f(M)=1.15\ M_{\odot}$, suggesting that the bright ($V=2.5$) B2IV star may host a dark, compact object companion. We use VLTI/GRAVITY observations to definitively test this possibility by directly resolving the binary. We detect a main sequence B star companion and rule out the compact object scenario. By combining the relative astrometric orbit and archival radial velocities, we report an updated precise characterization of the orbit (period $P=116.795\pm 0.002$ d, eccentricity $e=0.1764\pm 0.0004$, inclination $i=74.04\pm 0.01^{\circ}$) and estimate the masses of the B stars. Using the original Hipparcos parallax measurement $\varpi=6.05\pm 0.48$ mas, we find $M_{1}=10^{+4}_{-2}\ M_{\odot}$ and $M_{2}=6.9\pm 1.0\ M_{\odot}$. The uncertainties on the masses are primarily driven by the uncertain parallax, which we find is likely biased by the orbital motion. We use an archival UVES spectrum and MIST evolutionary tracks to refine our mass estimates. Finally, we discuss how interferometry and high-contrast imaging may be used to characterize other candidate star+compact object binaries, including those that will be discovered with Gaia DR4, as part of a larger effort to uncover the hidden population of black holes in the Milky Way.
@@ I Introduction
The majority of the observed black holes and neutron stars in the Milky Way have been detected in X-ray binary systems (Corral-Santana et al., 2016). While compact object masses can sometimes be measured dynamically in these systems, either by inferring the orbital inclination from the accretion disk properties (e.g., Torres et al., 2020) or by modeling ellipsoidal modulations (e.g., Orosz et al., 2001), many of the transient BH candidates have no mass measurements (e.g., Russell et al., 2022). Since X-ray binaries are expected to undergo at least one episode of binary interaction and mass transfer (Kalogera and Webbink, 1996; Tauris and van den Heuvel, 2006), the observed compact object mass distribution from X-ray binaries may present a biased view of the intrinsic distribution.
@@
Outside of the Galaxy, gravitational waves have been used to detect >150>150 compact object binary mergers (The LIGO Scientific Collaboration et al., 2025a). While these targets provide powerful constraints on the mass distribution of black holes (The LIGO Scientific Collaboration et al., 2025b), these rare extragalactic systems are also the products of complex binary interactions that result in a coalescence. The majority of the observed gravitational wave sources are also more massive than typical X-ray binaries, since more massive black hole mergers have a larger characteristic strain.
@@
In recent years, there has been an increased effort to uncover the hidden population of dormant black holes in the Milky Way. Isolated black holes can only be detected via microlensing (Lam et al., 2022; Lam and Lu, 2023). While there is currently only one strong candidate, the Vera C. Rubin Legacy Survey of Space and Time and the Roman Space Telescope are expected to detect $\gtrsim 30$ isolated black holes (Abrams et al., 2025; Lam et al., 2023). Multiple observational methods can be used to detect and characterize black holes in wide binaries with luminous companions. Three black holes have been detected using Gaia astrometry (El-Badry et al., 2023b, a; Chakrabarti et al., 2023; Tanikawa et al., 2023), including Gaia BH3, a red giant star in the Milky Way halo with a $33\ M_{\odot}$ black hole in an 11.9 year binary orbit (Gaia Collaboration et al., 2024). Spectroscopic surveys have also been applied to search for candidates (e.g., Giesers et al., 2018; Shenar et al., 2022; Mahy et al., 2022), but the unknown orbital inclination means that only a minimum black hole mass can be measured. At short orbital periods ($P\lesssim 10$ days for main sequence companions), ellipsoidal modulations can be used to measure the orbital inclination (Morris and Naftilan, 1993), but no strong candidates have been discovered starting from photometric variability (Rowan et al., 2021; Green et al., 2025).
@@
Regardless of whether a black hole candidate is detected with astrometry, spectroscopy, or photometry, confirming a black hole effectively means ruling out any type of luminous companion. For some systems, like the $33\ M_{\odot}$ black hole in Gaia BH3, the spectral energy distribution can easily exclude a massive main sequence star or an equal-mass binary of two $16.5\ M_{\odot}$ stars. For lower-mass black holes, spectral disentangling can help to distinguish compact objects from false positives, such as rapidly rotating subgiants (e.g., El-Badry et al., 2022; Jayasinghe et al., 2022). However, if we can directly resolve a candidate system with high-contrast imaging or interferometry, we can detect or rule out luminous companions with a single observation (e.g., Frost et al., 2022).
@@
If we observe a stellar binary, the direct imaging observations will resolve the two stars and help constrain the three-dimensional orbit. If we instead resolve a non-interacting black hole binary, we expect to observe a single point source. The combination of RVs and astrometry can therefore be used to unambiguously detect and characterize non-interacting black hole binary systems. In Rowan et al. (2025), we performed a pilot study to explore the feasibility of this method using SHARK-VIS (Pedichini et al., 2022, 2024) on the Large Binocular Telescope (Hill et al., 2010). We ruled out black holes for three of the four systems observed, and showed that two of the binaries selected are likely hierarchical triples. Since the SHARK-VIS observations only require $\sim 10$ min of on-sky observing time, direct imaging can be used to rapidly vet black hole candidates in long period orbits more efficiently than high-resolution spectroscopic follow-up.
@@
Here, we use VLTI/GRAVITY to observe $\kappa$ Velorum (hereafter $\kappa$ Vel), a bright ($V=2.5$), single-lined spectroscopic binary with an RV orbit that indicates a large, massive companion. Section §II describes how we selected this target as a potential black hole candidate. Section §III reports the VLTI/GRAVITY observations and the describes the data reduction. We unambiguously detect a second star in the GRAVITY observations and rule out a black hole. We then characterize the stellar binary by simultaneously fitting the archival RVs with the relative astrometry in Section §IV, model an archival high resolution spectrum (Section §V), and make comparisons to evolutionary tracks of massive stars (Section §VI). Finally, we discuss the prospects for observing other high-mass function single-lined spectroscopic binaries with interferometry and direct imaging in Section §VII.
II Target Selection
@@
We selected targets following the procedure outlined in Rowan et al. (2025). In short, we started from two large catalogs of spectroscopic binaries, the Ninth Catalog of Spectroscopic Binary Orbits (Pourbaix et al., 2004, hereafter SB9) and single-lined spectroscopic binaries from the Gaia DR3 non-single stars catalog (Gaia Collaboration et al., 2023; Gosset et al., 2025, hereafter Gaia SB1). We used the astrophysical parameter estimates from StarHorse (Anders et al., 2022) to estimate the mass of the photometric primaries, and determined the minimum companion masses and maximum projected orbital separations based on the archival RV orbits. Figure 1 shows the maximum projected orbital separations and minimum companion mass measurements.
@@
We identified $\kappa$ Vel (HD 81188, HIP 45941) as having a large minimum companion mass and an angular separation accessible with VLTI/GRAVITY. With an apparent magnitude of $V=2.5$, $\kappa$ Vel is one of the brightest stars in the Southern sky and has been targeted as part of a number of surveys, but the secondary has not previously been characterized. $\kappa$ Vel has been most commonly observed as part of interstellar medium absorption line studies (e.g., Crawford, 1991; Fruscione et al., 1994; Jenniskens et al., 1997; Crawford, 2002; Smith et al., 2013; Cox et al., 2017). $\kappa$ Vel was first identified as a binary candidate by Wright (1905) and the orbit was characterized by Curtis (1907) as part of the D.O. Mills Expedition from Lick Observatory to the Southern Hemisphere. Five additional measurements taken by Buscombe and Morris (1960) confirmed the RV variability and orbit. Table 1 lists the archival measurements. We assume an RV uncertainty of 5 km/s for the Curtis (1907) measurements since measurement errors were not reported. Chini et al. (2012) also measured four RVs of $\kappa$ Vel as part of a spectroscopic surveys of massive stars. They classify the system as an SB1, but the individual RV measurements or spectra are not available. $\kappa$ Vel is included in the SB9 catalog with an orbital period of $P=116.65$ days, $K_{1}=46.5$ km/s, and a binary mass function $f(M)=1.15\ M_{\odot}$, where
@@
$M_{1}$ is the photometric primary mass, $M_{2}$ is the companion mass, $i$ is the orbital inclination, and $e$ is the orbital eccentricity. The photometric primary is spectroscopically characterized as a B2IV star (Levato, 1975), which roughly corresponds to a primary mass of $\sim 7$$10\ M_{\odot}$ (Pecaut and Mamajek, 2013). Based on the RV orbit, even if the primary is $7\ M_{\odot}$ and the orbit is edge-on, the companion must be $5.7\ M_{\odot}$ (Figure 2). Because of the large minimum companion mass, Trimble and Thorne (1969) included $\kappa$ Vel in their catalog of binaries that could host compact object companions. The companion could be another massive star, a compact object, or an inner binary of two lower mass stars. The maximum projected orbital separation is >7.0>7.0 mas (Figure 1), making this a suitable target for VLTI direct imaging observations to test these possibilities.
Figure 1: Minimum companion masses, $M_{2,\rm{min}}$, and maximum projected angular separation for binaries in the Gaia SB1s and SB9 catalogs. $\kappa$ Vel is marked in blue. The vertical line marks the projected angular separation resolvable with VLTI/GRAVITY.
Table 1: Archival RV measurements of $\kappa$ Vel from Curtis (1907) and Buscombe and Morris (1960). Since Curtis (1907) do not report RV uncertainties, we estimate a conservative RV uncertainty of 5 km/s. We also include our RV measurement from the UVES spectrum. We assign an RV uncertainty on the UVES spectrum of 22 km/s to account for the unknown systematic RV offset between the archival measurements and the UVES measurement.
@@
III VLTI/Gravity Observations
@@
Table 2: VLTI/GRAVITY observation log. The dates are given for the start of the night. The calibrator uniform-disk (UD) diameters are estimated with the JMMC SearchCal tool (Chelli et al., 2016).
@@
III.1 Astrometry model-fitting
The visibility and closure phase data shows strong sinusoidal modulations that are indicative of a close binary system. In order to quantify the binary star parameters, we fit geometric models with the PMOIRED modeling software (Mérand, 2022). The free parameters are the relative separation vector of the two stars measured from the primary to the secondary (dRA, dDEC) and their $K$-band flux ratio ($\epsilon=f_{2}/f_{1}$). While we resolve the two stars, we do not resolve their stellar disks, so we fix the disk diameters to $0.1$ mas and use a uniform-disk model. For the error estimation, we bootstrap over the individual measurements that were taken during each epoch. The derived astrometry is reported in Table 3.
Table 3: Astrometry for $\kappa$ Vel. The measurement uncertainties are estimated by bootstrapping over the individual measurements taken during each epoch (Section §III.1). $\rho$ is the correlation coefficient between the dRA and dDEC measurements and $f_{2}/f_{1}$ is the flux ratio in the $K$-band.
@@ IV Orbit Fitting
We simultaneously model the RVs and relative astrometry with Markov Chain Monte Carlo (MCMC) with emcee (Foreman-Mackey et al., 2013). We sample over six orbital parameters: the period, $P$, the periastron time, $T_{0}$, the orbital inclination, $i$, the longitude of the ascending node, $\Omega$, the eccentricity, $e$, and the longitude of periastron, $\omega$. We parameterize the latter two as $\sqrt{e}\cos\omega$ and $\sqrt{e}\sin\omega$ for computational efficiency. The angular size of the astrometric orbit is the ratio of the orbital semi-major axis to the distance, $a_{0}=a/d$. We sample over the velocity-semi amplitude of the photometric primary, $K_{1}$, and the system velocity, $\gamma$. We include an offset $\delta_{\rm{RV}}$ between the Curtis (1907) and Buscombe and Morris (1960) RVs to account for systematic differences between the instruments. We use a Gaussian prior on $\delta_{\rm{RV}}$ centered on zero with $\sigma=1.5$ km/s. Since there is only one UVES spectrum, we do not fit for the potential RV offset between the Curtis (1907) RVs and the UVES spectrograph, and instead choose to inflate the measured UVES RV uncertainty (see Section §V).
Finally, we include two additional terms to account for potentially underestimated uncertainties. The stellar jitter term, $s$, accounts for stellar variability and underestimated RV uncertainties (Price-Whelan et al., 2017). To account for underestimated astrometric uncertainties, we also include a multiplicative term, $s_{\rm{ast}}$, that scales the bootstrapped astrometric uncertainties.
@@
where $\text{RV}_{\text{pred}}(t_{i})$ is calculated using the standard Keplerian orbit equations (Binnendijk, 1960). The astrometric orbit term is
@@
$G$ $=a_{0}\left(-\sin\Omega\sin\omega+\cos i\,\cos\Omega\cos\omega\right).$
We use 20 walkers and run the chains for 500,000 iterations. We visually inspect the walker distributions and discard the first 100,000 iterations as burn-in. We find that the Gelman-Rubin statistic is $\hat{R}<1.01$ and the effective sample size is >10,000>10{,}000 for all parameters, suggesting that the chains have converged. Figure 3 shows the resulting orbit model and Figure 4 shows the corner plot of the MCMC posteriors and random samples of the orbit model compared to the VLTI astrometry observations. Table 4 reports the median and $1\sigma$ uncertainties of the posteriors.
Given a parallax, $\varpi$, and associated uncertainty, $\sigma_{\varpi}$, the binary masses $M_{1}$ and $M_{2}$ can be solved for using Kepler’s third law. Figure 5 shows the mass posteriors using the original Hipparcos parallax $\varpi=6.05\pm 0.48$ mas (ESA, 1997). The revised Hipparcos parallax, which primarily includes improvements to attitude reconstruction, is $\varpi=5.7\pm 0.3$ mas (van Leeuwen, 2007). The revised Hipparcos astrometry is a “stochastic” solution, rather than the standard five-parameter solution, which includes an additional “cosmic noise” astrometric noise term. Unfortunately, although $\kappa$ Vel has been observed by Gaia, since it is $G=2.5$ mag, it is likely saturated (Fabricius et al., 2021) and has no reported parallax.
@@
$\alpha=a\left(\frac{q}{1+q}-\frac{\epsilon}{1+\epsilon}\right).$ (8)
@@
We must estimate the flux ratio of the system in the Hipparcos $H_{p}$ filter to determine the parallax bias. For a given mass pair of $M_{1}$ and $M_{2}$ and metallicity, we interpolate the MIST evolutionary tracks (Dotter, 2016; Choi et al., 2016) onto the same age grid, calculate the $K$-band flux as a function of stellar age, and use the observed flux ratio to estimate the system age. We use the $H_{p}$-band flux ratio at this system age to calculate $\alpha\sin i$. We find $\alpha\sin i\approx 0.22$, which indicates that the parallax can be biased by the binary motion by as much as 22%. Since the mass ratio depends on the parallax, this fractional bias does increase with increasing distance, but only at the percent level.
Figure 3: Top: RV orbit model and residuals. The red lines show random samples from the MCMC posteriors. Bottom left: VLTI/GRAVITY astrometry measurements (black points) and random samples for the astrometric orbit model (blue). Bottom right: residuals for the astrometric orbit model. The solid errorbars give the uncertainty derived from the PMOIRED model, and the dashed errorbars report the scaled uncertainty $\sigma_{i}^{\prime}=s_{\text{ast}}\sigma_{i}$.
@@
V UVES Spectra
$\kappa$ Vel was previously observed with VLT UVES as part of the ESO Diffuse Interstellar Bands Large Exploration Survey (EDIBLES, Cox et al., 2017, 194.C-0833). $\kappa$ Vel was observed on 2017-04-18 in all four UVES cross-dispersers (CDs). The Phase 3 data product includes the combined spectra for 20 individual exposures in CD1 and CD2 and for 14 and 15 exposures in CD3 and CD4, respectively. While the purpose of the EDIBLES program was to characterize diffuse interstellar bands (DIBs), we can also use the UVES $\kappa$ Vel spectrum to characterize the photometric primary of the binary. We start by continuum normalizing the CD2 spectrum with a cubic spline, which covers a wavelength range of 370–500nm. We mask the Balmer lines during continuum normalization. Our analysis only uses the CD2 observations given the density of spectral features and lack of telluric lines.